OpenStudy (anonymous):

3squared + (7 - 2) · 4 - 6 over 3

OpenStudy (anonymous):

is it similar to this? \[3^2 + (7-2) \times (4-6) \over 3\]

OpenStudy (anonymous):

yeah

OpenStudy (theviper):

Then it will be done as :- \[\Large{\frac{9+5\times-2}{3}}\]\[\Large{\frac{-1}{3}}\]

OpenStudy (anonymous):

possible answer choices are 25, 54, 27, 28

OpenStudy (anonymous):

24*

OpenStudy (theviper):

Choices are wrong :P

OpenStudy (anonymous):

lol cant be

OpenStudy (theviper):

ok I can also be wrong so I should call one more helper :) @hartnn :)

hartnn (hartnn):

the question is \[3^2+(7-2)*4-\frac{6}{3}\] \[=9+20-2=27\]

OpenStudy (anonymous):

ahhh that is why -_-

OpenStudy (theviper):

yes the question is wrong :P

OpenStudy (theviper):

that was given by @JusticeArchetype :P

OpenStudy (anonymous):

and He agreed that it is similar LOL

hartnn (hartnn):

and incorrectly confirmed by @ash.ferrante

OpenStudy (theviper):

yes he should sure confirm :) It was @ash.ferrante fault .

OpenStudy (anonymous):

(5+2)sqaured-40+3

hartnn (hartnn):

7^2=49 49-40=9 9+3=12